OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±(5/4)x

mathslover (mathslover):

The transverse axis is parallel to y axis . (h,k) is the centre (say) - (0,0)

mathslover (mathslover):

Now, a = distance from the center to the vertices along the transverse axis b = distance from the center to the endpoints of the conjugate axis c = distance from the center to the foci along the transverse axis it is given that, a= 10 here. The relationship established between a, b and c is : \(\bf 10^2 + b^2 = c^2\) Given, slope of asympotes is \(\bf{\pm \cfrac{5}{4} }\) a = is a multiple of 5 b = is a multiple of 4.

mathslover (mathslover):

As, a = 10. So, by symmetry , b = 8

mathslover (mathslover):

\(\bf{10^2 + 8^2 = c^2}\) \(\bf{100+64=c^2}\) \(\bf{164 = c^2}\) \(\bf{c = 2\sqrt{41}}\)

mathslover (mathslover):

Can you write equation of hyperbola now?